How many gallons of a 70% antifreeze solution must be mixed with 70 gallons of 10% antifreeze to get a mixture that is 60% antifreeze?
Answer by stanbon(75887)

How many gallons of a 70% antifreeze solution must be mixed with 70 gallons of 10% antifreeze to get a mixture that is 60% antifreeze?
---------------
Equation:
alch + alch = alch
0.70x + 0.10*70 = 0.60(x+70)
-----
70x + 10*70 = 60x + 60*70
-----
10x = 50*70
x = 5*70
x = 350 gallons (amt. of 70% solution to add)
----------------
Cheers,
Stan H.
------------